A stone is thrown upward from the top of a building at an angle of 30.0° to the


horizontal and with an initial speed of 20.0 m/s. The point of release is 45.0 m above the


ground.


a. How long does it take to hit the ground?


b. Find the stone’s speed impact.


c. Find the horizontal range of the stone.

1
Expert's answer
2022-02-11T08:47:26-0500

a. We have to solve the equation h = h0 + V0yt - ½gt² when h = 0 m as the ball hits the ground, h0 is the height of the building, and the velocity will be V0y = V0•sin(30°)=(20 m/s)(½)=10 m/s. Using that information, the quadratic equation that has to be solved for the time t is "45+10t-4.9035t^2= 0." Solving that we find the time as t = 4.216 s.


b. The velocity will be the combination of the velocity for the x and y axis. First:

"V_{0x}= V_{0}\cos(30°)=(20\frac{m}{s})(\frac{\sqrt{3}}{2})=17.32\frac{m}{s}
\\ V_{0y}= V_{0}\sin(30°)=(20\frac{m}{s})(\frac{1}{2})=10\frac{m}{s}"

Then we have to consider that the new velocity for the y axis has to be calculated while the other velocity remains constant:

"V_y=V_{0y}-gt
\\ V_y=[10-(9.807)(4.216)]\frac{m}{s}=-31.346\frac{m}{s}"

The total velocity will be "V=\sqrt{V_y^2+V_x^2}
\\ V=\sqrt{(-31.346\frac{m}{s})^2+(17.32\frac{m}{s})^2}
\\ V= 35.81\frac{m}{s}"


c. The distance that the stone travels through the x axis will be equal to "d = V_{0x}t= (17.32\frac{m}{s})(4.216\,s)=73.02\,m"


In conclusion, the stone travels a time of 4.216 s, hits the ground with a velocity of 35.81 m/s, and hits the ground at a distance of 73.02 m from the place where the stone was thrown.